# AdvRepair: Provable Repair of Adversarial Attack

### v1

@@ Abstract
Deep neural networks (DNNs) are increasingly deployed in safety-critical domains, but their vulnerability to adversarial attacks poses serious safety risks. Existing neuron-level methods using limited data lack efficacy in fixing adversaries due to the inherent complexity of adversarial attack mechanisms, while adversarial training, leveraging a large number of adversarial samples to enhance robustness, lacks provability. In this paper, we propose AdvRepair, a novel approach for provable repair of adversarial attacks using limited data. By utilizing formal verification, AdvRepair constructs patch modules that, when integrated with the original network, deliver provable and specialized repairs within the robustness neighborhood. Additionally, our approach incorporates a heuristic mechanism for assigning patch modules, allowing this defense against adversarial attacks to generalize to other inputs. AdvRepair demonstrates superior efficiency, scalability and repair success rate. Different from existing DNN repair methods, our repair can generalize to general inputs, thereby improving the robustness of the neural network globally, which indicates a significant breakthrough in the generalization capability of AdvRepair.
1 Introduction
@@
In the last few years, deep neural networks (DNNs) have made significant advancements in various domains, including computer vision [3], natural language processing [16], and speech recognition [71]. However, the adoption of DNNs in safety-critical domains has been slow due to concerns regarding their dependability. One major concern is the vulnerability of DNNs to adversarial attacks [43, 14], where adversaries can manipulate input data in a way that is imperceptible to humans but can cause the model to make incorrect decisions. This vulnerability poses serious safety risks in applications such as autonomous vehicles [8] and medical diagnosis [68]. Therefore, it is crucial to develop and deploy DNN models with safety and reliability requirements in mind to fully realize their potential in safety-critical domains.
The repair of DNNs involves rectifying errors in the network by modifying its architecture or parameters. These errors can arise from various sources, including training errors [10], adversarial attacks [26, 13], backdoor attacks [40], and distribution drift [78, 62]. Regular and efficient repair during DNN deployment is crucial to mitigate the risks associated with these errors and ensure the performance and reliability of the network in practical applications. Traditional machine learning approaches for DNN repair include techniques such as adversarial training [24, 23, 65], input sanitization, fine-tuning, transfer learning [15, 81], and data augmentation [82, 51, 42]. These methods rely on a large amount of data and may exhibit suboptimal performance in situations with limited or low-quality data. Neuron-level fault localization and repair methods [61] aim to improve the accuracy and reliability of DNN repair by identifying and correcting errors at the individual neuron level. By leveraging these neuron-level techniques, researchers and practitioners can achieve more precise and reliable DNN repair with limited data.
At present, the efficacy of neuron-level repair methods utilizing only a limited number of samples for robust defense against adversarial attacks remains limited. These methods demonstrate effectiveness primarily in the context of backdoor attacks or individual safety properties. This limitation arises from the inherent complexity of adversarial attack mechanisms, which often exhibit a higher degree of intricacy compared to backdoor attacks. Moreover, diverse adversarial attacks targeting distinct samples correspond to a unique dense mixture of features [1], rendering it arduous to generalize parameter adjustments at the neuron level based on a small sample set. Furthermore, while techniques such as adversarial training leverage a large number of adversarial samples to bolster the robustness of neural networks, their efficacy lacks provability. Consequently, neural networks repaired using such techniques continue to harbor latent vulnerabilities in terms of robustness.
Our proposed methodology aims to address the issue of adversarial attacks by leveraging a limited set of adversarial samples. Specifically, our approach focuses on rectifying these erroneous samples while simultaneously maximizing accuracy and ensuring local robustness on these samples. Moreover, we aim to achieve this local robustness in a manner that generalizes across the entire dataset. To achieve these objectives, our method integrates an external indicator, a hard-coded characteristic function, which facilitates the allocation of patch modules in conjunction with the original network. This indicator identifies inputs within the local neighborhood of a specific sample and assigns the appropriate patch modules for repair. Additionally, for inputs outside the local neighborhood, a heuristic allocation mechanism is employed to enable generalization across the entire dataset.
Each patch module constitutes a fully connected neural network structure. The core concept of our approach lies in the utilization of reachability analysis through a linear approximation for the provable training of patch modules. Specifically, we employ the verification method DeepPoly [58] to establish a linear approximation of the output neurons. This approximation is leveraged to determine the distance between the target behavior and the current behavior, serving as the loss function. The patch modules are then trained to minimize this distance. Once the loss function reaches zero, the patch modules provide provable repairs for adversarial attacks within the perturbation region. By employing an external indicator to identify faulty inputs, adversarial attacks within the same perturbation region can be repaired using the same patch module. Inputs outside the perturbation range of limited adversarial samples are assigned by the indicator, leveraging a heuristic allocation mechanism, thereby ensuring the generalization of the repair. Through the comprehensive framework of our method, we effectively enhance the resilience of the network against adversarial attacks while preserving accuracy. By leveraging reachability analysis and patch modules, we provide provable repairs for adversarial samples, thereby improving the robustness of the network.
The contributions of this paper can be summarized as follows:
  • We propose a novel approach, named AdvRepair, specifically designed for provable repair of adversarial attacks using limited data. This work represents a significant breakthrough in achieving provable repair for adversarial robustness. Our method introduces patch modules that are seamlessly integrated into the original network architecture. By combining a loss function derived from formal methods, we train these patch modules to provide provable repairs for local robustness. Through a heuristic allocation mechanism, our approach successfully achieves the generalization of local robustness repair.
@@
  • To address the efficiency challenges associated with formal verification in large-scale DNNs, we leverage patch modules to perform repairs at the feature space of the networks. This strategy enables our method to scale effectively across various network architectures, ensuring efficient and practical implementation.
@@
  • We conduct a comprehensive evaluation of AdvRepair on three diverse datasets and multiple DNN architectures. Through extensive comparisons with state-of-the-art repair and adversarial training approaches, our method consistently demonstrates superior efficiency and scalability. Moreover, it exhibits remarkable generalization capabilities to handle general inputs, thereby significantly enhancing the overall robustness of the network. This advancement in generalization capabilities represents a substantial contribution of our method.
Overall, our proposed approach, AdvRepair, offers a pioneering solution for provable repair of adversarial attacks using limited data. It not only addresses the challenges posed by formal verification in large-scale DNNs but also achieves remarkable generalization capabilities, leading to improved network robustness.
2 Preliminary
In this section we recall some basic notations of DNN repair.
@@ 2 Preliminary
A deep neural network is a function ${N}\colon\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}^{n_{L}}$ that maps an input $x\in\mathbb{R}^{n_{0}}$ to an output $y\in\mathbb{R}^{n_{L}}$. We usually visualize a DNN $N$ as a sequence of $L$ layers, where the $i$th layer contains $n_{i}$ neurons each representing a real variable. Between two adjacent layers is typically a composition of an affine function and a non-linear activation function, and the DNN $N$ is the composition of the functions between layers. In many applications, DNNs are serving for classification tasks. In such a classification DNN ${N}\colon\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}^{n_{L}}$, every output dimension corresponds to a classification label, and the one with the maximum output value is the classification results that the DNN $N$ gives, i.e., $C_{N}(x)=\arg\max_{1\leq i\leq n_{L}}N(x)_{i}$, where $N(x)_{i}$ is the $i$th entry of the vector $N(x)$.
@@
The notion of safety properties pertains to assertions that guarantee the absence of undesirable behavior. Within the context of DNNs, a safety property demands that a DNN operates correctly within a specified input range. The local robustness of a DNN refers to its ability to maintain stable and consistent predictions in the vicinity of its training data points, even in the presence of small perturbations or variations in the input. Here a neighborhood of an input $x_{0}$ is usually defined as a closed ball $B(x_{0},r):=\{x\in\mathbb{R}^{n_{0}}\mid\|x-x_{0}\|_{\infty}\leq r\}$, where $\|\cdot\|_{\infty}$ is the $L_{\infty}$-norm, and $r>0$ is the radius. Formally, a local robustness property of a classification DNN $N$ requires that for any input $x$ in a given neighborhood $B(x_{0},r)$ of an input $x_{0}$, its classification should always be consistent with $x_{0}$, i.e., $\forall x\in B(x_{0},r),C_{N}(x)=C_{N}(x_{0})$. We denote this local robustness property as $(N,B(x_{0},r))$; it is obviously a safety property. An input $x^{*}\in B(x_{0},r)$ is an adversarial sample of the DNN $N$, if $C_{N}(x^{*})\neq C_{N}(x_{0})$, i.e., $x^{*}$ is a counterexample which violates the local robustness property $(N,B(x_{0},r))$.
@@
Given a DNN $N$ and a set of adversarial samples $\{x_{i}^{*}\}_{i=1}^{n}$, where $n$ can be significantly smaller than the size of the dataset, and each $x_{i}^{*}$ is obtained by an adversarial attack on an input $x_{i}$ with a given radius $r$, we need to construct a DNN $F$ which is locally robust on every $B(x_{i},r)$ while the accuracy is maintained and local robustness of other inputs is potentially improved.
3 Methodology
@@
Here the first objective is the fundamental requirement showing that the repair is successful on the given adversarial samples $\{x_{i}^{*}\}_{i=1}^{n}$. The second objective aims to ensure the local generalizability of the repair, whereby the absence of adversarial examples is guaranteed within the vicinity of the generated adversarial examples, while the third objective focuses on the global generalizability of the repair, wherein it simultaneously improves the robustness against all potential inputs in the dataset without significantly compromising the accuracy.
@@
The current neural network repair methods under limited data conditions primarily involve parameter adjustments at the neuron level. While these methods demonstrate high efficacy against straightforward backdoor attacks, they face significant challenges in effectively addressing complex adversarial attacks. Even if the erroneous behaviors of $\{x_{i}^{*}\}_{i=1}^{n}$ are successfully repaired, such repairs lack the ability to generalize either locally or globally. Inspired by Reassure [22], we propose a patch-based repair method. A patch refers to a specific modification or alteration made to a software system or codebase; it is a discrete set of changes applied to fix a bug, enhance functionality, or address security vulnerabilities. Patch-based DNN repair involves the integration of an external indicator designed to identify buggy inputs, followed by the application of specialized patch modules to repair input sets which have similar behaviors. Each indicator here corresponds precisely to a robustness neighborhood $B(x_{i},r)$ that requires repair, and we can leverage verification tools to ensure that the repair within each robustness neighborhood is provable. Furthermore, for other inputs in the dataset, we can heuristically match patches that maximize their robustness, thereby endowing this repair approach with global generalization. In this section, we propose a patch-based repair method named AdvRepair to defend from adversarial attacks with limited data.
3.1 Structure of the repaired DNN
@@
The main produce of the repaired network is depicted in Fig. 1. For an input $x\in\mathbb{R}^{n_{0}}$, the role of the indicator is to select an appropriate patch that, when applied, yields the sum of the output of the patch module and the output of the original DNN for $x$. This sum represents the output obtained after the repair process.
Figure 1: Framework of AdvRepair
The definition of an indicator function is standard:
Definition 1
@@
Let $\mathcal{C}=(X_{1},\ldots,X_{m})$ be a finite sequence of input properties. The indicator function $\mathcal{I}_{\mathcal{C}}\colon\mathbb{R}^{n_{0}}\to\{0,1\}^{m}$ outputs in the $j$th entry as
@@
$\mathcal{I}_{\mathcal{C}}(x)_{j}=\left\{\begin{array}[]{ll}1,&\text{ if }x\models X_{j},\\ 0,&\text{ otherwise},\end{array}\right.$
where $j=1,\ldots,m$.
@@
Upon classification by the indicator, a set of patch modules is deployed to perform the repair. Here each patch module is specifically tailored to address an input set with the same local robustness property. The implementation of a patch module is a fully connected neural network in this work. For an input $x$ that is not within the input set of any indicator, i.e., $x\notin\bigcup_{i=1}^{n}B(x_{i},r)$, it cannot be identified by the indicators, but there is still potential adversarial attack in $B(x,r)$. In this situation, we heuristically allocate a patch module to this input to defend from adversarial attacks. Formally, the structure of the repaired DNN is as follows:
Definition 2
@@
  • and $\tau\colon\mathbb{R}^{n_{0}}\setminus\bigcup_{i=1}^{n}B(x_{i},r)\to 2^{\{1,\ldots,m\}}$ is a patch allocation function.
The semantics of the repaired DNN $F=(N,\mathcal{C},\mathcal{P},\tau)$ is a function
@@
$x$$\mapsto\left\{\begin{array}[]{ll}\mathcal{I}_{\mathcal{C}}(x)^{\mathrm{T}}\mathcal{P}(x)+N(x),&\text{ if }x\in\bigcup_{i=1}^{n}B(x_{i},r),\\ \sum_{j\in\tau(x)}P_{j}(x)+N(x),&\text{ otherwise},\end{array}\right.$
@@
Generally there are three ways to determine the number of the patch modules and define the indicator:
@@
  • Property-guided. The property-guided way sets $\mathcal{C}_{\mathrm{P}}=\{B(x_{i},r)\mid i=1,\ldots,n\}$, i.e., the indicator judges which robustness region the input belongs to, and the corresponding patch module defends against adversarial attacks in this specific region.
  • Untargeted label. A simple label-guided way takes every classification label as an input property. i.e., $\mathcal{C}_{\mathrm{L}}=\{1,\ldots,n_{L}\}$, where $x\models\ell$ iff $C_{N}(x)=\ell$. Assuming that $x$ is correctly classified by $N$, the patch module to defend against adversarial attacks in $B(x,r)$ is determined by $C_{N}(x)$, and it works for fixing potential adversarial attacks of all the robustness regions where the ground truth is $C_{N}(x)$.
  • Targeted label. We consider the set of label pairs $\mathcal{C}_{\mathrm{L}}=\{(\ell_{0},\ell_{1})\in\{1,\ldots,n_{L}\}^{2}\mid\ell_{0}\neq\ell_{1}\}$, where $x\models(\ell_{0},\ell_{1})$ iff $N(x)_{\ell_{0}}>N(x)_{\ell_{1}}>N(x)_{\ell}$ for all $\ell\neq\ell_{0},\ell_{1}$, i.e., the indicator determines which label $\ell_{1}$ is the most likely when attacking the ground truth $\ell_{0}$, assuming that $x$ is correctly classified by $N$, and the corresponding patch module fixes adversarial attacks of this specific pattern.
The property-guided way has advantages when the number $n$ of the given adversarial samples is small, while the label-guided ways work well in the situation when the number $n_{L}$ of the output labels is small. For the label-guided ways, the patch allocation function $\tau$ can be simply defined as the corresponding indicator, since this indicator can naturally generalize to other input. In the following, we present our approach in the property-guided way, and it can be easily adapted to the label-guided ways.
3.2 Training the patch modules

### v2

# Patch Synthesis for Property Repair of Deep Neural Networks

@@ Abstract
Deep neural networks (DNNs) are prone to various dependability issues, such as adversarial attacks, which hinder their adoption in safety-critical domains. Recently, NN repair techniques have been proposed to address these issues while preserving original performance by locating and modifying guilty neurons and their parameters. However, existing repair approaches are often limited to specific data sets and do not provide theoretical guarantees for the effectiveness of the repairs. To address these limitations, we introduce PatchPro, a novel patch-based approach for property-level repair of DNNs, focusing on local robustness. The key idea behind PatchPro is to construct patch modules that, when integrated with the original network, provide specialized repairs for all samples within the robustness neighborhood while maintaining the network’s original performance. Our method incorporates formal verification and a heuristic mechanism for allocating patch modules, enabling it to defend against adversarial attacks and generalize to other inputs. PatchPro demonstrates superior efficiency, scalability, and repair success rates compared to existing DNN repair methods, i.e., realizing provable property-level repair for 100% cases across multiple high-dimensional datasets.
I Introduction
@@
In recent years, deep neural networks (DNNs) have achieved significant advancements in various domains, including computer vision [1], natural language processing [2], and speech recognition [3]. Despite these advancements, the adoption of DNNs in safety-critical domains has been slow due to concerns regarding their dependability. A major concern is the vulnerability of DNNs to adversarial attacks [4, 5], where adversaries can manipulate input data in ways that are imperceptible to humans but can cause the model to make incorrect decisions. This vulnerability poses serious safety risks in applications such as autonomous vehicles [6] and medical diagnosis [7]. Therefore, it is crucial to regularly update the DNN to mitigate the risks associated with these errors and ensure the reliability of the network in practical applications.
DNN repair techniques [8, 9, 10] have been proposed to address the problem involving rectifying errors in the network by modifying its architecture or parameters. Compared to traditional methods such as adversarial training [11, 12, 13], input sanitization, fine-tuning, transfer learning [14, 15], and data augmentation [16, 17, 18], neuron-level fault localization and repair methods [8] offer a more targeted approach by identifying and correcting errors at the individual neuron level while not affecting the overall network performance.
Despite significant advancements in this field, several limitations remain. First, neuron-level repair techniques, which rely on a limited number of samples, often struggle to provide robust defense against adversarial attacks due to the inherent complexity of these attacks compared to simpler threats like backdoor attacks. Adversarial attacks involve intricate mixtures of features [19], making it difficult to generalize parameter adjustments from a small dataset. Second, existing repair methods typically focus on specific data for repair and fail to generalize to the property level (e.g., local robustness), which limits their effectiveness in addressing a broad range of adversarial scenarios. This motivates us to explore property repair of DNNs, specifically to fix the safety properties that neural networks violate in certain input regions. Third, while provable repair methods such as PRDNN [20], REASSURE [21], and APRNN [22] can conduct property-based error correction, they achieve this by ensuring that outputs meet constraints on the vertices of the input region as a polyhedron. However, the number of vertices increases exponentially with data dimensionality, which makes these methods inefficient for high-dimensional data.
In this work, we aim to address these limitations by proposing a novel patch-based method to achieve provable repair on the property level. Specifically, we focus on correcting potential adversarial samples in the infinite set of high-dimensional points within a certain neighborhood of these error samples, i.e., satisfying local robustness property. To achieve this, our key idea is to use formal verification to help construct a separate patch module (in the form of a fully connected neural network structure) for each neighborhood outside of the original network. Such a patch-based method allows us not only to ensure that the infinite set of points within the error sample’s neighborhood repaired but also to maintain the performance of the original network unaffected. To construct such a patch, we utilize reachability analysis through linear relaxation for the provable training of these patch modules. Specifically, we use the verification method DeepPoly [23] to create a linear relaxation of the output neurons. This approximation is employed to determine the distance between the targeted behavior and the current behavior, serving as the loss function. The patch modules are then trained to minimize this distance. Once the loss function reaches zero, the patch modules will offer provable repairs for adversarial attacks within the perturbation region.
To ensure that each patch module addresses specific neighborhoods, our approach employs an external indicator that identifies inputs within a particular sample’s local neighborhood and assigns the appropriate patch modules for repair. This indicator allows the same patch module to effectively repair adversarial attacks within the same perturbation region. For adversarial samples outside the repaired property, the indicator utilizes a heuristic allocation mechanism to assign suitable patch modules, thereby enhancing the generalization of the repair. This comprehensive framework not only improves the network’s resilience against adversarial attacks but also maintains its accuracy. By leveraging reachability analysis and dedicated patch modules, we provide provable repairs for adversarial samples, significantly boosting the network’s robustness. Additionally, to extend this local robustness across the entire dataset, we integrate the external indicator with the original network, ensuring that all samples, including those beyond the initially repaired property, benefit from heuristic patch module allocation. This approach effectively combines local and global repair strategies to enhance overall network robustness.
We summarize our contributions as follows:
  • We introduce PatchPro, a novel approach that integrates a loss function derived from formal verification to train patch modules within the original network. This method effectively repairs high-dimensional infinite point sets within the polyhedron neighborhood of adversarial samples while preserving the network’s performance. The approach also includes a heuristic allocation mechanism, which ensures the generalization of local robustness repair by seamlessly integrating patch modules into the original network architecture.
@@
  • To tackle the efficiency challenges of formal verification in large-scale DNNs, we utilize patch modules to perform repairs in the feature space of the networks. This strategy enables our method to scale effectively across various network architectures, ensuring both efficient and practical implementation.
@@
  • We thoroughly evaluate PatchPro on three diverse datasets and multiple DNN architectures. Through extensive comparisons with state-of-the-art repair and adversarial training techniques, our method consistently demonstrates superior efficiency, scalability, and generalization capabilities. It shows significant improvement in handling general inputs, thereby greatly enhancing the overall robustness of the network.
II Preliminary
@@
In this section we recall some basic notations of DNN repair. A deep neural network is a function ${N}\colon\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}^{n_{L}}$ that maps an input $x\in\mathbb{R}^{n_{0}}$ to an output $y\in\mathbb{R}^{n_{L}}$. We usually visualize a DNN $N$ as a sequence of $L$ layers, where the $i$th layer contains $n_{i}$ neurons each representing a real variable. Between two adjacent layers is typically a composition of an affine function and a non-linear activation function, and the DNN $N$ is the composition of the functions between layers. In many applications, DNNs are serving for classification tasks. In such a classification DNN ${N}\colon\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}^{n_{L}}$, every output dimension corresponds to a classification label, and the one with the maximum output value is the classification result that the DNN $N$ gives, i.e., $C_{N}(x)=\arg\max_{1\leq i\leq n_{L}}N(x)_{i}$, where $N(x)_{i}$ is the $i$th entry of the vector $N(x)$.
The notion of safety properties pertains to assertions that guarantee the absence of undesirable behavior. Within the context of DNNs, a safety property demands that a DNN operates correctly within a specified input range.
Definition 1.
A safety property is a triple $({N},{X},{Q})$, where $N$ is a DNN, ${X}\subseteq\mathbb{R}^{d_{\text{in }}}$ and ${Q}\subseteq\mathbb{R}^{d_{\text{out }}}$ are the subset of input and output spaces of the neural network ${N}$. The property $({N},{X},{Q})$ is satisfied if and only if ${N}(x)\in{Q}$ for all $x\in{X}$.
A local robustness property of a classification DNN $N$ requires that for any input $x$ in a given neighborhood $B(x_{0},r)$ of an input $x_{0}$, its classification should always be consistent with $x_{0}$, where a neighborhood of an input $x_{0}$ is usually defined as a closed ball $B(x_{0},r):=\{x\in\mathbb{R}^{n_{0}}\mid\|x-x_{0}\|_{\infty}\leq r\}$, $\|\cdot\|_{\infty}$ is the $L_{\infty}$-norm, and $r>0$ is the radius. Formally, it can be defined as follows:
@@
The local robustness of a DNN refers to its ability to maintain stable and consistent predictions in the vicinity of its in-distribution data points, even in the presence of small perturbations or variations in the input. Here a neighborhood of an input $x_{0}$ is usually defined as a closed ball $B(x_{0},r):=\{x\in\mathbb{R}^{n_{0}}\mid\|x-x_{0}\|_{\infty}\leq r\}$, where $\|\cdot\|_{\infty}$ is the $L_{\infty}$-norm, and $r>0$ is the radius. Formally, a local robustness property of a classification DNN $N$ requires that for any input $x$ in a given neighborhood $B(x_{0},r)$ of an input $x_{0}$, its classification should always be consistent with $x_{0}$, i.e., $\forall x\in B(x_{0},r),C_{N}(x)=C_{N}(x_{0})$. We denote this local robustness property as $(N,B(x_{0},r))$, i.e. $(N,B(x_{0},r))=(N,B(x_{0},r),\{y\in\mathbb{R}^{n_{L}}\mid y_{i}<y_{C_{N}(x_{0})},i=1,2,\dots,n_{L}\})$, and thus it is a safety property.
@@
Given a DNN $N$ and a set of adversarial samples $\{x_{i}^{*}\}_{i=1}^{n}$, where $n$ can be significantly smaller than the size of the training set, and each $x_{i}^{*}$ is obtained by an adversarial attack on an input $x_{i}$ with a given radius $r$, we need to construct a DNN $F$ which is locally robust on every $B(x_{i},r)$ while the accuracy is maintained and local robustness of other inputs is potentially improved.
III Methodology
@@
In this work, we propose a patch-based repair method. A patch refers to a specific modification or alteration made to a software system or codebase; it is a discrete set of changes applied to fix a bug, enhance functionality, or address security vulnerabilities. Patch-based DNN repair involves the integration of an external indicator designed to identify buggy inputs, followed by the application of specialized patch modules to repair input sets which have similar behaviors. Each indicator here corresponds precisely to a robustness neighborhood $B(x_{i},r)$ that requires repair, and we can leverage verification tools to ensure that the repair within each robustness neighborhood is provable. Furthermore, for other inputs in the dataset, we can heuristically match patches that maximize their robustness, thereby endowing this repair approach with global generalization. In this section, we propose a patch-based repair method named PatchPro to defend from adversarial attacks with limited data.
III-A Structure of the repaired DNN
@@
The main produce of the repaired network is shown in Fig. 1. For an input $x\in\mathbb{R}^{n_{0}}$, the role of the indicator is to select the appropriate patches that, when applied, yields the sum of the outputs of the patch modules and the output of the original DNN for $x$. This sum represents the output obtained after the repair process. The indicator function is defined as:
Figure 1: The architecture of a DNN repaired by PatchPro. It contains multiple patch networks for each input properties. Each of them is enabled or disabled according to the allocation signal “1” or “0” determined by the indicator. The blue lines highlights the indicator’s workflow. The final output is the sum of the outputs of all the enabled patches and the original network.
Definition 2.
@@
Let $\mathcal{C}=(X_{1},\ldots,X_{m})$ be a finite sequence of input properties. The indicator function $\mathcal{I}_{\mathcal{C}}\colon\mathbb{R}^{n_{0}}\to\{0,1\}^{m}$ outputs in the $j$th entry, where $j\in\{1,\dots,m\}$, as
@@
$\mathcal{I}_{\mathcal{C}}(x)_{j}=\left\{\begin{array}[]{ll}1,&\text{ if }x\models X_{j},\\ 0,&\text{ otherwise}.\end{array}\right.$
Typically an input property $X_{i}$ is a subset of $\mathbb{R}^{n_{0}}$, and we define $x\models X_{i}$ iff $x\in X_{i}$.
@@
Upon classification by the indicator, a set of patch modules is deployed to perform the repair. Here each patch module is specifically tailored to address an input set with the same local robustness property. The implementation of a patch module is a fully connected neural network in this work. For an input $x$ that does not satisfy any input properties, i.e., $x\notin\bigcup_{i=1}^{n}B(x_{i},r)$, there is no existing specific patch module, but it may still suffer from adversarial attacks within its neighborhood. In this situation, we heuristically allocate some patch modules to this input to defend from adversarial attacks. Formally, the structure of the repaired DNN is as follows:
Definition 3.
@@
  • and $\tau\colon\mathbb{R}^{n_{0}}\setminus\bigcup_{i=1}^{n}X_{i}\to 2^{\{1,\ldots,m\}}$ is a patch allocation function.
The semantics of the repaired DNN $F$ is a function:
@@
$x$$\mapsto\left\{\begin{array}[]{ll}N(x)+\mathcal{I}_{\mathcal{C}}(x)^{\mathrm{T}}\mathcal{P}(x),&\text{ if }x\in\bigcup_{i=1}^{n}X_{i},\\ N(x)+\sum_{j\in\tau(x)}P_{j}(x),&\text{ otherwise},\end{array}\right.$
@@
In this work, we set $\mathcal{C}=\{B(x_{i},r)\mid i=1,\ldots,n\}$, i.e., the indicator judges which robustness region the input belongs to, and the corresponding patch module defends against adversarial attacks in this specific region.
III-B Training the patch modules